OpenStudy (anonymous):

If ABC,and DEF are similar and, BC = 5 feet, EF = 11 feet, and FD = 3 feet, find AC ? as a fraction

OpenStudy (anonymous):

this makes no sense

OpenStudy (anonymous):

no

OpenStudy (anonymous):

no wut???

OpenStudy (anonymous):

Can you help me turtleluver ?

Directrix (directrix):

If two triangles are similar, then the ratio of corresponding sides are in proportion. That gives: BC/EF = AC/DF 5/11 = (AC)/3 11(AC) = 15 AC = ? @jayy32